OpenStudy (anonymous):

What is the equation of the parabola with a directrix at y = -4 and focus (0, 4)?

OpenStudy (mertsj):



OpenStudy (anonymous):

 x^2 = 4py x^2 = 4*4y x^2 = 16y y = 1/16(x^2)